OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by the equation below. y = 2x^2 + 2x - 4

OpenStudy (yrelhan4):

find the values of x for y=0..

OpenStudy (irishboy123):

divide by 2 x^2+x-2 = 0 (x+2)(x-1)=0